# Large Language Models for Secure Code Assessment: A Multi-Language Empirical Study

### v1

@@
Abstract
Most vulnerability detection studies focus on datasets of vulnerabilities in C/C++ code, offering limited language diversity. Thus, the effectiveness of deep learning methods, including large language models (LLMs), in detecting software vulnerabilities beyond these languages is still largely unexplored.
In this paper, we evaluate the effectiveness of LLMs in detecting and classifying Common Weakness Enumerations (CWE) using different prompt and role strategies. Our experimental study targets six state-of-the-art pre-trained LLMs (GPT-3.5-Turbo, GPT-4 Turbo, GPT-4o, CodeLLama-7B, CodeLLama-13B, and Gemini 1.5 Pro) and five programming languages: Python, C, C++, Java, and JavaScript. We compiled a multi-language vulnerability dataset from different sources, to ensure representativeness. Our results showed that GPT-4o achieves the highest vulnerability detection and CWE classification scores using a few-shot setting.
Aside from the quantitative results of our study, we developed a library called CodeGuardian integrated with VSCode which enables developers to perform LLM-assisted real-time vulnerability analysis in real-world security scenarios. We have evaluated CodeGuardian with a user study involving 22 developers from the industry. Our study showed that, by using CodeGuardian, developers are more accurate and faster at detecting vulnerabilities.
I Introduction
A software vulnerability is a defect that could allow an attacker to gain control of a software system, steal or manipulate sensitive data, install a backdoor, or plant other types of malware [1, 2, 3]. Software vulnerabilities vary in nature and level of exploitability. Not all vulnerabilities allow attackers to cause the same types of harm, and not all vulnerabilities create equally severe risks. However, all vulnerabilities pose at least some level of risk to the applications they impact, as well as the environments that host those applications.
Software vulnerabilities may have different root causes, spanning from the way the software is designed, problems with the software’s source code, poor management of data, or access control settings within the application. In this work, we focus on classes of vulnerabilities related to bugs in the source code. The U.S. Department of Homeland Security reports that 90% of security incidents are caused by exploiting security flaws and defects in the code [4]. A recent survey [5] found that over 55% of developers struggle to identify code vulnerabilities, indicating a lack of secure coding knowledge. Due to the increasing frequency and severity of cybersecurity incidents, which negatively impact end-users, businesses, and critical infrastructure [6], identifying security vulnerabilities within source code has become essential.
Researchers have proposed methods for the automatic detection and repair of software vulnerabilities [7, 8, 9, 10, 11, 12, 13]. For instance, program analysis-based techniques have been widely employed [14]. However, such rule-based techniques are affected by high false positive rates and low effectiveness in handling a large number of diverse types of vulnerabilities [12]. In addition, they often suffer from long response times [13], which makes them less suitable for continuous integration/development environments. Recently, Deep Learning (DL) techniques have enabled significant progress in vulnerability detection in the source code [15, 16, 17]. However, most DL models treat the code as linear sequences or textual data, which limits the finding of the semantic relationship within the source code and demands the development of more robust and efficient methods.
To address these limitations, researchers have started exploring the potential of Large Language Models (LLMs) that are pre-trained on a large corpus of data and use advanced natural language processing capabilities based on the transformer architecture [18]. LLMs have shown great results across various natural language and software engineering tasks [18, 11], including automated vulnerability detection and repair [9, 10, 11]. Existing studies on vulnerability detection are executed against datasets that include vulnerabilities pertaining to only one or two programming languages. Instances are the vulnerability-fixing commit dataset by Zhou et al. [19] (C/C++), by Li et al. [20] (Java), BigVul [21] (C/C++), SVEN [22] (Python and C/C++), Devign [23] (C), D2A [24] (C/C++). Consequently, the capabilities of LLMs across diverse programming languages remain largely unexplored. Furthermore, these studies are purely quantitative, and the potential of LLMs to assist developers in identifying vulnerabilities has not yet been investigated.
To fill these research gaps, this paper investigates the effectiveness of LLMs in identifying vulnerable codes across different programming languages. To the best of our knowledge, this is the first study that reports an extensive evaluation of six state-of-the-art pre-trained LLMs (GPT-3.5-Turbo, GPT-4 Turbo, GPT-4o, CodeLLama-7B, CodeLLama-13B, and Gemini 1.5 Pro) in identifying the vulnerabilities pertaining to five programming languages, namely Python, C, C++, Java, and JavaScript. Our study focuses on the top 25 Common Weakness Enumeration (CWE) classes that represent the most dangerous software weaknesses [25], for which we compiled a dataset of vulnerable snippets from existing reference datasets. Our study targets vulnerability detection and CWE classification using zero-shot and few-shot approaches.
Furthermore, to investigate the industrial applicability of LLMs in software vulnerability detection and debugging, we have developed a VSCode extension called CodeGuardian. It integrates the LLM analyzed by our quantitative study to allow developers to scan code dynamically while editing, providing instant feedback and enhancing secure coding practices directly within their workflow. We have evaluated CodeGuardian with a user study involving six tasks across the five programming languages and 22 participants with software engineering experience in the industry and varying levels of expertise in software security. Our qualitative analysis revealed that by using CodeGuardian, developers were twice as accurate at CWE classification and 60% faster, thanks to the LLM-powered engine.
Our paper makes the following contributions:
Dataset. A dataset of more than 370 manually validated vulnerabilities, pertaining to five programming languages, which can support the evaluation of automated techniques in security vulnerability detection and classification.
Evaluation. An empirical study concerning the capabilities of six pre-trained LLMs at detecting and classifying software vulnerabilities in five programming languages.
CodeGuardian. A VSCode extension that allows developers to use LLMs to scan code for vulnerabilities while editing.
User Study. A qualitative analysis of CodeGuardian with a user study of 22 developers from our industrial partner.
II Empirical Study
II-A Research Questions
@@
RQ2 (CWE classification): How effective are LLMs in classifying CWEs across multiple languages? How does effective vary across languages? How does effectiveness vary among different languages?
The first research question evaluates how accurately LLMs perform binary classification, determining whether a code snippet is vulnerable or not. The second research question evaluates the accuracy of LLMs in classifying CWEs by assigning the correct CWE ID to a given code snippet.
II-B Dataset Design
The representativeness of datasets is crucial for effective vulnerability detection and classification [26, 27]. Existing works adopted datasets with limited language diversity, focusing primarily on C/C++ [19, 20, 21, 22, 23, 24].
In contrast, in this work, we consider software vulnerabilities of five popular programming languages, namely Python, C, C++, Java, and JavaScript. We compiled a dataset that integrates three existing vulnerability datasets: CVEFixes, CWE-snippets, and the JavaScript Vulnerability DataSet (JVD). Table I shows the statistics of these datasets, including the number of languages, CWEs, and files. CVEFixes is a fix-commit vulnerability dataset [28]. CWE-snippets is a dataset created in collaboration with an industrial partner for this study. It consists of numerous vulnerability snippets across multiple programming languages. JVD is a vulnerability dataset for JavaScript code [29]. From each dataset, we selected the top 25 CWE most dangerous software weaknesses [25] for each of the five selected programming languages.
Table I: Vulnerabilities benchmarks evaluated in this study.
@@ II-B Dataset Design
      JVD [29]       1       N/A       12,125
The maximum number of snippets per class was set to three across all the languages due to a limited number of vulnerable code samples available in the reference datasets for certain CWE classes and programming languages. For example, CWE-787 (Out-of-bounds Write) is more commonly associated with memory-unsafe languages like C and C++, while it is less prevalent in memory-safe languages such as Python and Java. Consequently, we prioritized constructing a balanced representation of vulnerabilities across all languages in our dataset. As a result, we aimed to construct a balanced representation of vulnerabilities across all languages in our dataset. In total, we retained 378 snippets, ensuring an equal number of vulnerable and non-vulnerable snippets for each programming language ($n_{Python}=38$, $n_{Java}=42$, $n_{C++}=36$, $n_{C}=44$, and $n_{JavaScript}=29$).
Table II: LLMs evaluated for vulnerability detection and classification.
@@
GPT-3.5-Turbo [30] $>$ 100B Decoder 16,385 4,096
GPT-4 Turbo [31] $>$ 100B Decoder 128K 4,096
GPT-4o [32] $>$ 100B Decoder 128K 4,096
CodeLLama-7B [33] 7 B Decoder 100K 4,096
CodeLLama-13B [33] 13 B Decoder 100K 4,096
Gemini 1.5 Pro [34] $>$ 100B Decoder 1M 8,192
Table III: System and user prompts: VD-Sys1, VD-Sys2, VD-User1 and VD-User2 are used for vulnerability detection (RQ1). CWE-Sys1 and CWE-Sys2 are used for system prompts 1 and 2 for CWE Classification. CWE-UserZ (zero-shot) and CWE-UserF (few-shot) are user prompts used for CWE classification (RQ2).
@@
VD-Sys1 System Description You are an AI binary vulnerability classifier that identifies whether the provided code is vulnerable or not vulnerable. You should respond with either only ’vulnerable’ or ’not vulnerable’. N/A
VD-Sys2 System Description You are an experienced developer who knows the security vulnerability very well. N/A
VD-User1 User Description for VD Classify the following code in vulnerable or not vulnerable. Output either only ’vulnerable’ or ’not vulnerable’. +: vulnerable
-: non-vulnerable
VD-User2 User Description for VD Now you need to identify whether a code contains a vulnerability or not. If has any potential vulnerability, output: ’vulnerable’. Otherwise, output: ’not vulnerable’.
@@
+: vulnerable
-: non-vulnerable
CWE-Sys1 System Description for CWE You are an AI vulnerability classifier that identifies CWE of a sent code, if that code has a vulnerability.
@@
CWE-Sys2 System Description for CWE Classification You are an experienced developer who knows the security vulnerability very well. N/A
CWE-UserZ User Description for CWE (zero-shot) Classify the following code in CWE a category. Output only one CWE tag in lowercase letters. Do not provide the full name of the CWE. Respond with ’non-vul’ if you think the code is not vulnerable. CWE ID
CWE-UserF User Description for CWE Classification with the few-shot setting Classify the following code in a CWE category. The code should have a vulnerability that corresponds to one of the top-25 CWEs below. Output only one CWE tag in lowercase letters. Do not provide the full name of the CWE. Respond ’non-vul’ if you think the code is not vulnerable. Below is the list of the top 25 CWEs:
@@
II-C Large Language Models
We evaluate the effectiveness of using six state-of-the-art pre-trained LLMs. Table II shows the selected LLMs, including their parameter size (the number of parameters for GPT and Gemini 1.5 Pro models are not disclosed by the organizations, and therefore, we report an estimate), architecture, input, and output token sizes. Specifically, we considered GPT-3.5 Turbo, GPT-4 Turbo, and CodeLlama models available as of June 2024, and pertaining from different organizations such as OpenAI’s GPT-3.5-Turbo [30], GPT-4 Turbo [31] and GPT-4o [32], as well as Meta’s CodeLLama-7B [33] and CodeLLama-13B [33], and Google’s Gemini 1.5 Pro [34].
For the GPT models, we utilized the supporting company’s Azure OpenAI Studio. Concerning hosting CodeLLama-7B and CodeLLama-13B models, we employed Google Colab and the transformers library by HuggingFace [35]. For the Gemini 1.5 Pro model, we used the Vertex AI from the Google Cloud Platform. Lastly, to minimize the models’ creativity and obtain more focused, conservative, and consistent responses, we set the temperature parameter to $0.1$ for all LLMs, as done in previous studies [36].
Table IV: RQ1: Effectiveness comparison of six LLMs using four prompt configurations for CWE detection. The best scores are indicated in bold.
LLM / Prompt VD-Sys1 + VD-User1 VD-Sys1 + VD-User2 VD-Sys2 + VD-User1 VD-Sys2 + VD-User2
Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$

### v2

@@
Abstract.
Most vulnerability detection studies rely on datasets tied to specific programming languages and within-project settings, limiting language diversity and the assessment of model generalizability. Consequently, the effectiveness of deep learning methods, including large language models (LLMs), beyond such restricted scenarios remains largely unexplored. In this paper, we evaluate six state-of-the-art LLMs (GPT-3.5-Turbo, GPT-4 Turbo, GPT-4o, CodeLlama-7B, CodeLlama-13B, Gemini 1.5 Pro) on vulnerability detection and CWE classification across five languages (Python, C, C++, Java, JavaScript). We compiled a multi-language dataset from diverse sources and investigated different prompt and role strategies. Our results show that GPT-4o achieves the best performance, particularly in few-shot settings. To examine practical applicability, we developed CodeGuardian, a VSCode extension that enables real-time LLM-assisted vulnerability detection. In a user study with 22 professional developers, CodeGuardian doubled accuracy and halved task completion time compared to a control group.
Vulnerability Detection, Large Language Models, Security, LLMs
journalyear: 2026copyright: ccconference: 7th International Workshop on Deep Learning for Testing and Testing for Deep Learning (DeepTest ’26); April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 7th International Workshop on Deep Learning for Testing and Testing for Deep Learning (DeepTest ’26) (DeepTest ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3786154.3788584isbn: 979-8-4007-2386-5/2026/04ccs: Software and its engineering Software verification and validation
[…]
by
1. Introduction
Software vulnerabilities are defects that can enable attackers to steal data, gain control of systems, or insert malware (Shirey, 2007; Shoshitaishvili et al., 2015; Or-Meir et al., 2019). They arise from design flaws, insecure code, or poor data and access management. In this work, we focus on vulnerabilities rooted in source code bugs. According to the U.S. Department of Homeland Security, 90% of security incidents exploit such flaws (Cybersecurity and Infrastructure Security Agency (CISA), 2021). Yet, over half of developers report difficulties identifying vulnerabilities (GitLab, 2022), highlighting the need for effective detection tools.
Traditional program-analysis techniques (Pistoia et al., 2007; Pinconschi et al., 2021; Thapa et al., 2022; Alqarni and Azim, 2022; Chen et al., 2023b; Johnson et al., 2013; Rangnau et al., 2020) are limited by high false positives, narrow coverage, and long analysis times (Johnson et al., 2013; Rangnau et al., 2020). Deep learning approaches (Russell et al., 2018; Hanif et al., 2021; Dam et al., 2021; Humbatova et al., 2020) improved scalability but often treat code as text, failing to capture semantic structure (Humbatova et al., 2020; Riccio et al., 2020). Large Language Models (LLMs), pre-trained on massive code and text corpora (Vaswani et al., 2017), have recently shown promise for vulnerability detection (Thapa et al., 2022; Alqarni and Azim, 2022; Chen et al., 2023b). However, existing evaluations are limited to one or two languages (e.g., BigVul (Fan et al., 2020), SVEN (He and Vechev, 2023)), and LLM capabilities across diverse languages remain largely unexplored. Moreover, most studies remain purely quantitative, with little evidence on how LLMs can assist developers in practice.
This paper addresses these gaps by evaluating six state-of-the-art LLMs (GPT-3.5-Turbo, GPT-4 Turbo, GPT-4o, CodeLlama-7B, CodeLlama-13B, Gemini 1.5 Pro) on vulnerability detection and CWE classification across five languages (Python, C, C++, Java, JavaScript). We focus on the top 25 CWE classes (1) and compile a dataset of over 370 manually validated vulnerable snippets. Beyond our quantitative analysis, we investigate practical applicability through CodeGuardian, a VSCode extension that integrates LLMs into developers’ workflows. In a user study with 22 industry practitioners, CodeGuardian doubled accuracy in CWE classification and reduced task completion time by 60%. The contributions of this work are as follows:
Dataset. A curated dataset of 370+ vulnerabilities across five languages, supporting evaluation of detection and classification tools.
Evaluation. An empirical study of six LLMs on vulnerability detection and classification across multiple languages.
CodeGuardian. A VSCode extension enabling real-time vulnerability scanning, released in our replication package (rep-package, 2025).
User Study. A qualitative evaluation with 22 professional developers, showing significant improvements in speed and accuracy with CodeGuardian.
2. Empirical Study
2.1. Research Questions
@@
RQ2 (CWE classification): How effective are LLMs in classifying CWEs across multiple languages? How does effectiveness vary among different languages?
2.2. Dataset Design
The representativeness of datasets is crucial for effective vulnerability detection and classification (Ding et al., 2024; Croft et al., 2023). Most existing works adopted datasets with limited language diversity, focusing primarily on C/C++ (Zhou et al., 2024a; Li et al., 2023; Fan et al., 2020; He and Vechev, 2023; Zhou et al., 2019; Zheng et al., 2021). Although some researchers have attempted to address language diversity limitations with evaluating LLMs on three to seven programming languages (Shu et al., 2025; Atiiq et al., 2024; Zhang et al., 2025), they either focus solely on binary vulnerability detection without CWE classification (Shu et al., 2025; Atiiq et al., 2024). Additionally, most existing studies rely heavily on publicly available vulnerability data primarily from the National Vulnerability Database (NVD), potentially limiting the diversity and real-world applicability of their findings.
In contrast, in this work, we consider software vulnerabilities of five popular programming languages, namely Python, C, C++, Java, and JavaScript, which represent the top-ranked languages according to IEEE’s 2025 survey (IEEE Spectrum, 2025). Furthermore, we compiled a dataset that integrates three existing vulnerability datasets: CVEFixes, CWE-snippets, and the JavaScript Vulnerability DataSet (JVD). Table 1 shows the statistics of these datasets, including the number of languages, CWEs, and files. CVEFixes is a fix-commit vulnerability dataset (Bhandari et al., 2021). CWE-snippets is a dataset created in collaboration with an industrial partner for this study. It consists of numerous vulnerability snippets across multiple programming languages. JVD is a vulnerability dataset for JavaScript code (Ferenc et al., 2019). From each dataset, we selected the top 25 CWE most dangerous weaknesses (1) for each of the five selected programming languages.
Table 1. Vulnerabilities benchmarks evaluated in this study.
@@ 2.2. Dataset Design
JVD (Ferenc et al., 2019) 1 N/A 12,125
2.2.1. Dataset Construction and Validation
Given the critical importance of data quality in vulnerability detection research (Ding et al., 2024), we adopted a systematic approach for dataset construction that addresses known limitations in existing vulnerability datasets. Our methodology involves several key phases designed to ensure both representativeness and accuracy of the collected data.
The data collection process involved executing systematic queries against the CVEFixes and CWE-snippets databases to extract code snippets corresponding to specific CWE categories across all target programming languages. Since JVD was structured as a CSV file, we employed systematic filtering techniques to identify relevant JavaScript vulnerabilities. To maintain dataset balance and address the inherent scarcity of certain vulnerability types in specific languages, we set a maximum threshold of three snippets per CWE class across all languages. This constraint was particularly important given that certain vulnerabilities—such as CWE-787 (Out-of-bounds Write)—tend to be associated with memory-unsafe languages like C and C++, while remaining relatively uncommon in memory-safe languages such as Python and Java.
Furthermore, our methodology involved manual validation of all selected code snippets. Each snippet has been inspected to verify both the presence of the vulnerability and the associated CWE ID. This validation process was essential to mitigate the data quality issues that have been identified in existing vulnerability datasets (Ding et al., 2024).
For binary vulnerability detection tasks, we curated non-vulnerable code snippets from multiple sources within our reference datasets. These included fixed versions of vulnerable code from CVEFixes commit histories, solution implementations from the CWE-snippets dataset, and verified secure code samples from JVD. The selection process for non-vulnerable snippets required careful manual inspection to ensure they represented genuinely secure implementations rather than simply alternative vulnerable patterns. It is important to note that the majority of non-vulnerable snippets in our dataset are not direct fixed versions of the corresponding vulnerable snippets. Based on our manual inspection, only a limited number of snippet pairs existed where the non-vulnerable code was the direct fixed version of a specific vulnerable snippet (one pair in C++ and two pairs in C). The use of independent non-vulnerable samples would also support model generalizability by ensuring exposure to diverse secure coding patterns rather than specific fix implementations.
Our final dataset contains 378 manually validated code snippets, balanced across vulnerable and non-vulnerable samples for each language (Table 2). It includes Python (38/38), Java (42/42), C++ (36/36), C (44/44), and JavaScript (29/29). Most samples were drawn from the CWE-snippets and CVEFixes datasets, while JVD provided the non-vulnerable JavaScript cases. To account for the hierarchical nature of CWE, predictions matching a child CWE (e.g., CWE-78) for a snippet labeled with its parent (e.g., CWE-77) were considered correct, with the parent assigned as the final class.
Table 2. Number of code snippets (vul. / non-vul.).
Language CVE fixes CWE snippets JSV Total
C 15 / 7 29 / 37 0 / 0 44 / 44
C++ 12 / 15 24 / 21 0 / 0 36 / 36
Java 16 / 3 26 / 39 0 / 0 42 / 42
JavaScript 15 / 4 14 / 2 0 / 23 29 / 29
Python 22 / 38 16 / 0 0 / 0 38 / 38
Total 80 / 67 109 / 99 0 / 23 189 / 189
Table 3. LLMs evaluated in this work.
@@ 2.2.1. Dataset Construction and Validation
GPT-3.5-Turbo (OpenAI, 2024a) >> 100B Decoder 16,385 4,096
GPT-4 Turbo (OpenAI, 2024b) >> 100B Decoder 128K 4,096
GPT-4o (OpenAI, 2024c) >> 100B Decoder 128K 4,096
CodeLLama-7B (rozière2024codellamaopenfoundation) 7B Decoder 100K 4,096
CodeLLama-13B (rozière2024codellamaopenfoundation) 13B Decoder 100K 4,096
Gemini 1.5 Pro (Team, 2024) >> 100B Decoder 1M 8,192
Table 4. System and user prompts: Sys1, VD-Sys2, VD-User1 and VD-User2 are used for vulnerability detection (RQ1). CWE-Sys1 and CWE-Sys2 are used for system prompts 1 and 2 for CWE Classification. CWE-UserZ (zero-shot) and CWE-UserF (few-shot) are user prompts used for CWE classification (RQ2).
@@
V-Sys1 System Description You are an AI binary vulnerability classifier that identifies whether the provided code is vulnerable or not vulnerable. You should respond with either only ’vulnerable’ or ’not vulnerable’. N/A
V-Sys2 System Description You are an experienced developer who knows the security vulnerability very well. N/A
U1 User Description for VD Classify the following code in vulnerable or not vulnerable. Output either only ’vulnerable’ or ’not vulnerable’. vulnerable
or
non-vulnerable
U2 User Description for VD Now you need to identify whether a code contains a vulnerability or not. If has any potential vulnerability, output: ’vulnerable’. Otherwise, output: ’not vulnerable’.
@@
vulnerable
or
non-vulnerable
C-Sys1 System Description for CWE You are an AI vulnerability classifier that identifies CWE of a sent code, if that code has a vulnerability.
@@
C-Sys2 System Description for CWE Classification You are an experienced developer who knows the security vulnerability very well. N/A
UZ User Description for CWE (zero-shot) Classify the following code in CWE a category. Output only one CWE tag in lowercase letters. Do not provide the full name of the CWE. Respond with ’non-vul’ if you think the code is not vulnerable. CWE ID
UF User Description for CWE Classification with the few-shot setting Classify the following code in a CWE category. The code should have a vulnerability that corresponds to one of the top-25 CWEs below. Output only one CWE tag in lowercase letters. Do not provide the full name of the CWE. Respond ’non-vul’ if you think the code is not vulnerable. Below is the list of the top 25 CWEs:
@@
2.3. Large Language Models
We evaluate the effectiveness of using six state-of-the-art pre-trained LLMs. Table 3 shows the selected LLMs, including their parameter size (the number of parameters for GPT and Gemini 1.5 Pro models are not disclosed by the organizations, and therefore, we report an estimate), architecture, input, and output token sizes. Specifically, we considered GPT-3.5 Turbo, GPT-4 Turbo, and CodeLlama models available as of June 2024, and pertaining from different organizations such as OpenAI’s GPT-3.5-Turbo (OpenAI, 2024a), GPT-4 Turbo (OpenAI, 2024b) and GPT-4o (OpenAI, 2024c), as well as Meta’s CodeLLama-7B (rozière2024codellamaopenfoundation) and CodeLLama-13B (rozière2024codellamaopenfoundation), and Google’s Gemini 1.5 Pro (Team, 2024).
For the GPT models, we utilized the supporting company’s Azure OpenAI Studio. Concerning CodeLLama-7B and CodeLLama-13B models, we employed Google Colab and the transformers library by HuggingFace (et al., 2020). For the Gemini 1.5 Pro model, we used the Vertex AI from the Google Cloud Platform. Lastly, to minimize the models’ creativity and obtain more focused, conservative, and consistent responses, we set the temperature parameter to $0.1$ for all LLMs, as done in previous studies (Chen et al., 2023a).
Table 5. RQ1: Effectiveness comparison of six LLMs using four prompt configurations for CWE detection.
LLM / Prompt V-Sys1–U1 V-Sys1–U2 C-Sys2–U1 C-Sys2–U2
Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$